Prove that the function f(x, y) = |xy| / xy is integrable on R= [−1,1]×[−1,1].
1
Expert's answer
2019-10-06T11:42:37-0400

Define function "f" on "A=\{0\}\times[-1,1]\cup[-1,1]\times\{0\}" as "f(t)=0" for all "t\in A" . Since "|f(t)|=1" for all "t\in[-1,1]^2\setminus A", we have that "|f(t)|\le 1" on "[-1,1]^2", that is "f" is bounded on "[-1,1]^2".

Next, area of "A" equals 0, and "f" is continuous on "[-1,1]^2\setminus A" , because "[-1,1]^2\setminus A" is disjoint union of "(0,1]^2", "(0,1]\times[-1,0)", "[-1,0)^2" and "[-1,0)\times(0,1]" , and "f" is continuous on every of these sets.

Since "[-1,1]^2" is Jordan mesurable, "f" is bounded and continuous almost everywhere on "[-1,1]^2", by Lebesgue's criterion for Riemann integrability we obtain that "f" is integrable on "[-1,1]^2" .